# AI-Driven Early Mental Health Screening with Limited Data: Analyzing Selfies of Pregnant Women

### v1

@@ 4.4 Hardware-software Setup
Hardware: Intel(R) Xeon(R) CPU @ 2.20GHz with 32GB of RAM, running Linux Ubuntu 22.04.4 LTS, and equipped with an NVIDIA L4 GPU with 24GB of memory. Software: The source code was written in Python 3.10, mostly using PyTorch 2.3 for model training and inference. The Sentence-BERT [33] is implemented in SentenceTransformers library [^1] https://sbert.net. The LLaVA-NeXT and Kosmos-2 models – llava-hf/llava-v1.6-mistral-7b-hf and microsoft/kosmos2-patch14-224, respectively, are available at the HuggingFace [^2] https://huggingface.co, while GPT-4o was accessed via the OpenAI API.
@@ 5.1 Comparative Evaluation
Table 1 shows the results of the comparative evaluation experiment. Overall, EfficientNetV2 outperformed the compared models among the CNNs for both FER2013 and RAF-DB pre-training, while GPT-4o yielded the best performance among the VLMs. Pre-training on RAF-DB improved most metrics for EfficientNetV2 and ResNet-18, although a notable decrease in performance was observed for ResNet-50 and VGG11. Notably, ResNet-18 yield better perfomance than ResNet-50, despite its reduced size. This suggests that larger models might require more data to achieve better performance.
@@ 5.2 Sensitivity Analysis
While Table 1 focused on the default FFNN model (VLM-based), this experiment investigated the alternative (single hidden layer) FFNN model (Equation (1)). More specifically, it was analyzed the impact of increasing the model complexity by varying the number of hidden units ($h=4,8,16,\ldots,256$). Figure 5 shows the results for this experiment. The dashed line in each chart represents the highest value for the respective metric as reported in Table 1. Notably, the F1-score achieved with GPT-4o and the default classification model (56.0%) – indicated by the dashed line in the third chart (from left to right) – demarcates an empirical upper bound for this metric.
@@ 5.3 Limitations and Challenges
Learning the relationship between facial expressions and self-reported PHQ-4 scores from single selfies can be challenging with small datasets. When using CNNs, prediction is mainly influenced by features around the mouth, as evidenced by the Grad-CAM [35] attention maps for EfficientNetV2 in Figures 6a and 6b. In Figure 6a, the slight smile led the model to classify a positive sample (PHQ-4 $\geq$ 6) as negative. However, facial cues around the eyes might suggest a posed (non-Duchenne) smile, which is not necessarily a sign of happiness or well-being. This issue could be mitigated by incorporating facial action units (AUs) into the learning process, enhancing attention guidance to other critical regions, as recently proposed in a study on basic emotions [15].

### v2

@@ 4.4 Hardware-software Setup
Hardware: Intel(R) Xeon(R) CPU @ 2.20GHz with 32GB of RAM, running Linux Ubuntu 22.04.4 LTS, and equipped with an NVIDIA L4 GPU with 24GB of memory. Software: The source code was written in Python 3.10, mostly using PyTorch 2.3 for model training and inference. The Sentence-BERT (Reimers and Gurevych, 2019) is implemented in SentenceTransformers library [^3] https://sbert.net. The LLaVA-NeXT and Kosmos-2 models – llava-hf/llava-v1.6-mistral-7b-hf and microsoft/kosmos2-patch14-224, respectively, are available at the HuggingFace [^4] https://huggingface.co, while GPT-4o was accessed via the OpenAI API.
5 Results and Discussion
@@ 5.1 Comparative Evaluation
Table 1 shows the results of the comparative evaluation experiment. Overall, EfficientNetV2 outperformed the compared models among the CNNs for both FER2013 and RAF-DB pre-training, while GPT-4o yielded the best performance among the VLMs. Pre-training on RAF-DB improved most metrics for EfficientNetV2 and ResNet-18, although a notable decrease in performance was observed for ResNet-50 and VGG11. Notably, ResNet-18 yielded better performance than ResNet-50, despite its reduced size. This suggests that larger models might require more data to achieve better performance.
@@ 5.2 Sensitivity Analysis
While Table 1 focused on the default FFNN model (VLM-based), this experiment investigated the alternative (single hidden layer) FFNN model (Equation (1)). More specifically, it was analyzed the impact of increasing the model complexity by varying the number of hidden units ($h=4,8,16,\ldots,256$). Figure 5 shows the results of this experiment. The dashed line in each chart represents the highest value for the respective metric as reported in Table 1. Notably, the F1-score achieved with GPT-4o and the default classification model (56.0%) – indicated by the dashed line in the third chart (from left to right) – demarcates an empirical upper bound for this metric.
@@ 5.3 Limitations and Challenges
Learning the relationship between facial expressions and self-reported PHQ-4 scores from single selfies can be challenging with small datasets. When using CNNs, prediction is mainly influenced by features around the mouth, as evidenced by the Grad-CAM (Selvaraju et al., 2017) attention maps for EfficientNetV2 in Figures 6a and 6b. In Figure 6a, the slight smile led the model to classify a positive sample (PHQ-4 $\geq$ 6) as negative. However, facial cues around the eyes might suggest a posed (non-Duchenne) smile, which is not necessarily a sign of happiness or well-being. This issue could be mitigated by incorporating facial action units (AUs) into the learning process, enhancing attention guidance to other critical regions, as recently proposed in a study on basic emotions (Belharbi et al., 2024).
@@
AI systems are notorious for showing errors and biases in different racial groups (Nazer et al., 2023). Despite the significance of this issue and its implications for fairness and equity, this sensitive topic was not explored in the present study.